mykenzie:

Isla flipped a coin 30 times. The coin landed heads up 9 times and tails up 21 times. Part A: Based on the results, what is the experimental probability of the coin landing heads up? Show your work. (5 points) Part B: What is the theoretical probability of the coin landing heads up? Show your work. (5 points)

Imagine:

A.)9/30=?

Kylo:

A) 9/30 or 3/10 B ) 9/30 so heads land up 9 times and the total flips are 30 so probability is 9/30=3/10

Imagine:

I think it needs to be a percent, does it? @mykenzie

mykenzie:

it doesn't matter

Kylo:

well hope i helped!

mykenzie:

you both did thxs